The number of people infected after t days is P = 200 + 5t - t2. Find the number of days until the number of people infected is 176.
Answer by sowmya(32)

The number of people infected after t days is P = 200 + 5t - t2. Find the number of days until the number of people infected is 176.
solution:
Let the number of people infected afetr t days be P
P = 200 + 5t - t^2
we haev to find t when P = 176
we have to solve for t
200+ 5t - t^2 = 176
t^2-5t-24=0
t^2-8t+3t-24=0
(t-3)(t-8)=0
Hence t=3 or t=8